Find the slope-intercept equation for the line with the indicated slope and y-intercept. Slope-4/9 and y-intercept(0,4)

Answer by Boreal(15235)

It is y=mx+b, where m is the slope and b is y when x=0.
here, m=-4/9 and b=4
y=-(4/9)x+4